7. Aniline, C6H5NH2, when reacted with picric acid gives a derivative with an absorptivity of
134 cm−1
g
−1 L at 359 nm. What would be the absorbance of a 1.00 × 10−4 M solution of
reacted aniline in a 1.00-cm cell?
1
Expert's answer
2021-08-16T02:46:05-0400

Solution:

The Beer–Lambert law can be used.

The Beer–Lambert law can be expressed as:

A = ε × l × C

where:

A is absorbance (no units)

ε is molar absorptivity (cm−1 M-1)

l is the optical path length (cm)

C is the concentration


Convert [cm−1 g−1 L] to [cm−1 M-1]:

The molar mass of aniline is 93.13 g mol−1.

Therefore,

(134 cm−1 g−1 L) × (93.13 g mol−1) = 12479.42 cm−1 L mol−1 = 12479.42 cm−1 M-1


Thus,

A = (12479.42 cm−1 M-1) × (1.00 cm) × (1.00 × 10−4 M) = 1.2479 = 1.25

A = 1.25


Answer: The absorbance (A) of a solution is 1.25